1. 13. This regression output is used in estimating the Fama- French 3 factor model. Use it to answer the following questions: R
    1. Use the info to create the model's equation.

Mkt-RF

SMB

HML

RF

S&P 500

NASDAQ

1.250

-0.940

-0.560

0.100

1.441

0.015

Find Expected return given the following market performance

c. What is SMB and MHL in this model?
Answer #1

Solution :

a) R ​= α​+β1​(Mkt-RF​) + β2​(SMB)​ + β3(​HML)

= -0.043 + 0.875(Mkt-RF) + 0.074(SMB) + (-0.087)(HML)

= -0.043 + 0.875(1.250) + 0.074(-0.940) + (-0.087)(-0.560)

= -0.043 + 1.0937 - 0.0695 + 0.0487

= 1.0299

The expected return from the given market performance will 1.0299

b) SMB and HML are the risk factor, which helps to predict Mkt-RF more significantly. Because of these two factors the alpha i.e. intercepts get disappear and it is no longer statistically significant at conventional levels.